# Many-body quantum catalysts for transforming between phases of matter

### v1

@@ VI Applications to state preparation
$|\mathrm{triv.}\rangle\otimes|\mathrm{triv.}\rangle\xrightarrow{I\otimes\mathcal{U}_{a}}|\mathrm{triv.}\rangle\otimes|a\rangle\xrightarrow{\mathcal{V}}|\psi\rangle\otimes|a\rangle$ (67)
@@
$|\psi\rangle=\mathcal{U}\mathcal{U}_{a}^{\dagger}\mathcal{U}^{\dagger}\mathcal{U}_{a}|\mathrm{triv.}\rangle$ (68)
@@
If $H_{a}(t)$ consists of symmetric terms then so will $\tilde{H}_{a}(t)$ since $\mathcal{U}_{\mathrm{SPT}}$ is a symmetric QCA. Furthermore, $H_{a}(t)$ and $\tilde{H}_{a}(t)$ will have the same spatial locality. So, if $|a\rangle$ can be prepared in time $\tau$ using a symmetric Hamiltonian evolution, then $|\psi\rangle$ can be prepared in time $2\tau$ using a symmetric Hamiltonian evolution with the same locality. If we instead write $\mathcal{U}_{a}$ as a symmetric quantum circuit, then $\mathcal{U}\mathcal{U}_{a}^{\dagger}\mathcal{U}^{\dagger}$ will be a symmetric quantum circuit with the same depth and locality (up to a constant).
@@ VI.1 Generating SPT phases with long-range interactions
$H=\sum_{i<j}h_{i,j}$ (71)
where $h_{i,j}$ acts on sites $i$ and $j$ and is bounded by $||h_{i,j}||<d_{i,j}^{-\alpha}$ where $d_{i,j}$ is the distance between $i$ and $j$ and $\alpha$ is some positive constant. The catalyst we consider is the GHZ state, which we have seen serves as a catalyst for SPT phases in all dimensions. Efficient protocols to prepare GHZ states using power-law interactions have been studied extensively due to their application in state-transfer, see Ref. Chen et al. (2023) for a recent review. Using the resulting GHZ state as a catalyst, all of these protocols can be immediately translated protocols to prepare SPT phases  [^1] It should be noted that the protocols used in optimal state-transfer protocols of Refs. Eldredge et al. (2017); Tran et al. (2021) do not use symmetric Hamiltonians. However, it is straightforward to modify these protocols such that they are symmetric, without changing their efficiency. See, as a demonstration, the linear-depth circuit in Ref. Chen et al. (2024) for creating a GHZ state..
In general, the optimal time to make a GHZ state using an $\alpha$-decaying Hamiltonian is a complicated function of $\alpha$. Here, we mention as one example that, for $\alpha<D$ where $D$ is the spatial dimension, the time to make a GHZ state is independent of system size. Therefore, SPT phases can be made in constant time using $\alpha$-decaying symmetric Hamiltonians for $\alpha<D$.
@@ VI.2 Generating SPT phases with local symmetric channels
$|a(\vec{s})\rangle=\prod_{i}\frac{\mathbb{1}+s_{i}Z_{i}Z_{i+2}}{2}|+\rangle^{\otimes N}$ (72)
@@
$|a(\vec{s})\rangle\langle a(\vec{s})|=\prod_{i=1}^{N-2}\left(\frac{\mathbb{1}+s_{i}Z_{i}Z_{i+2}}{2}\right)\left(\frac{\mathbb{1}+U_{e}}{2}\frac{\mathbb{1}+U_{o}}{2}\right)$ (73)
@@
Therefore, to prepare a cluster state, we can prepare a catalyst by performing symmetric local measurements on a product state and then applying the catalyzed symmetric FDQC described in Theorem 1. The catalyst that we prepare on average in the process is
$\rho=\sum_{\vec{s}}p_{\vec{s}}|a(\vec{s})\rangle\langle a(\vec{s})|=\frac{1}{2^{N}}(\mathbb{1}+U_{e})(\mathbb{1}+U_{o})$ (74)
which is exactly the mixed-state catalyst described in Sec. V. Therefore, the above process describes an operational “unravelling” what it means to use the mixed-state catalyst in practice.
@@ VII Discussion
In this work, we introduced the idea of a many-body quantum catalyst and demonstrated its richness and broad applicability in the context of SPT phases of matter. There are many avenues of future work, and we describe some here. First, it is worth noting that a majority of our results used a particular subclass of catalysts that we called $(G,\mathcal{U})$-symmetric catalysts. Can we find catalysts for SPT phases with new physical properties if we move away from this subclass? Or is it possible that all catalysts can be related in a particularly way to a $(G,\mathcal{U})$-symmetric catalyst? In general, it is an important problem to constrain potential catalysts at a more basic level than we have done here.
@@
Another future direction invovlving fermions is to make a connection to the notion of fragile topology Po et al. (2018), wherein a topologically nontrivial energy band can be trivialized by the addition of trivial degrees of freedom. This represents an analog of our results for non-interacting fermions. Whether ideas introduced in our paper, possibly in combination with ideas from fragile topology, can be employed to develop an understanding of many body catalysts for interacting fermionic systems would be an interesting topic for future work.
We have also focused on catalyzing short-range entangled phases of matter, and we suspect that catalysts will play a much different role when considering long-range entangled phases. As a preliminary comment, we remark that some of the aspects of our story are reminscent of the idea of foliated fracton phases of matter Shirley et al. (2018). Therein, certain 3D phases of matter called fracton phases have the property that an FDQC acting on the fracton state and a product state can turn that product state into a topologically ordered state. This process changes the 3D state by reducing its size in one direction, but preserves its phase of matter. So one can view a fracton phase as a catalyst for topologically ordered states, with the caveat that it can only be used finitely many times (assuming a finite system size).
In a similar vein, catalysts can also be used to transform between gapless phases of matter Scaffidi et al. (2017); Verresen et al. (2021). For example, it is shown in Ref. Verresen et al. (2021) that two gapless topological phases with distinct patterns of symmetry enrichment can be related by applying symmetric QCA corresponding to SPT entanglers. Since the catalysts we construct allow such QCA to be implemented on arbitrary states, they can also catalyze the transformation between different certain symmetry-enriched gapless phases.

### v2

@@ VI Applications to state preparation
$|\mathrm{triv.}\rangle\otimes|\mathrm{triv.}\rangle\xrightarrow{I\otimes\mathcal{U}_{a}}|\mathrm{triv.}\rangle\otimes|a\rangle\xrightarrow{\mathcal{V}}|\psi\rangle\otimes|a\rangle,$ (67)
@@
$|\psi\rangle=\mathcal{U}\mathcal{U}_{a}^{\dagger}\mathcal{U}^{\dagger}\mathcal{U}_{a}|\mathrm{triv.}\rangle.$ (68)
@@
where we have defined the modified Hamiltonian,
$\tilde{H}_{a}(t)=\mathcal{U}H_{a}(t)\mathcal{U}^{\dagger}.$ (71)
If $H_{a}(t)$ consists of symmetric terms then so will $\tilde{H}_{a}(t)$ since $\mathcal{U}_{\mathrm{SPT}}$ is a symmetric QCA. Furthermore, $H_{a}(t)$ and $\tilde{H}_{a}(t)$ will have the same spatial locality. Note that this only holds if $H_{a}$ is a symmetric Hamiltonian, which is where the symmetry of the catalyst comes in. So, if $|a\rangle$ can be prepared in time $\tau$ using a symmetric Hamiltonian evolution, then $|\psi\rangle$ can be prepared in time $2\tau$ using a symmetric Hamiltonian evolution with the same locality, without requiring ancillary degrees of freedom. If we instead write $\mathcal{U}_{a}$ as a symmetric quantum circuit, then $\mathcal{U}\mathcal{U}_{a}^{\dagger}\mathcal{U}^{\dagger}$ will be a symmetric quantum circuit with the same depth and locality (up to a constant).
@@ VI.1 Generating SPT phases with long-range interactions
$H=\sum_{i<j}h_{i,j},$ (72)
where $h_{i,j}$ acts on sites $i$ and $j$ and is bounded by $||h_{i,j}||<d_{i,j}^{-\alpha}$ where $d_{i,j}$ is the distance between $i$ and $j$ and $\alpha$ is some positive constant. The catalyst we consider is the GHZ state, which we have seen serves as a catalyst for SPT phases in all dimensions. Efficient protocols to prepare GHZ states using power-law interactions have been studied extensively due to their application in state-transfer, see Ref. Chen et al. (2023) for a recent review. Using the resulting GHZ state as a catalyst, all of these protocols can be immediately translated protocols to prepare SPT phases  [^1] It should be noted that the protocols used in optimal state-transfer protocols of Refs. Eldredge2017; Tran2021 do not use symmetric Hamiltonians. However, it is straightforward to modify these protocols such that they are symmetric, without changing their efficiency. See, as a demonstration, the linear-depth circuit in Ref. Chen et al. (2024) for creating a GHZ state..
In general, the optimal time to make a GHZ state using an $\alpha$-decaying Hamiltonian is a complicated function of $\alpha$. Here, we mention as one example that, for $\alpha<D$ where $D$ is the spatial dimension, the time to make a GHZ state is independent of system size. Therefore, SPT phases can be made in constant time using $\alpha$-decaying symmetric Hamiltonians for $\alpha<D$. To the best of our knowledge, this is the first construction of symmetric long-range Hamiltonians for preparing SPT phases.
@@ VI.2 Generating SPT phases with local symmetric channels
$|a(\vec{s})\rangle=\prod_{i}\frac{\mathbb{1}+s_{i}Z_{i}Z_{i+2}}{2}|+\rangle^{\otimes N},$ (73)
@@
$|a(\vec{s})\rangle\langle a(\vec{s})|=\prod_{i=1}^{N-2}\left(\frac{\mathbb{1}+s_{i}Z_{i}Z_{i+2}}{2}\right)\left(\frac{\mathbb{1}+U_{e}}{2}\frac{\mathbb{1}+U_{o}}{2}\right),$ (74)
@@
Therefore, to prepare a cluster state, we can prepare a catalyst by performing symmetric local measurements on a product state and then apply the catalyzed symmetric FDQC described in Theorem 1. The catalyst that we prepare on average in the process is
$\rho=\sum_{\vec{s}}p_{\vec{s}}|a(\vec{s})\rangle\langle a(\vec{s})|=\frac{1}{2^{N}}(\mathbb{1}+U_{e})(\mathbb{1}+U_{o}),$ (75)
which is exactly the mixed-state catalyst described in Sec. V. Therefore, the above process describes via an operational “unravelling” of what it means to use the mixed-state catalyst in practice. We stress that this mixed state is different from the pure GHZ state. In particular, it does not allow for long-range state transfer, so there is no contradiction that we are able to prepare it with a symmetric local channel.
@@ VII Discussion
In this work, we introduced the idea of a many-body quantum catalyst and demonstrated its richness and broad applicability in the context of SPT phases of matter. There are many avenues of future work, and we describe some here. First, it is worth noting that a majority of our results considered a particular class of catalysts that we called $(G,\mathcal{U})$-symmetric catalysts (Theorem 1), but we have not shown that all catalysts must belong to this class. Can we find catalysts for SPT phases with new physical properties if we move away from this class? Or is it possible that all catalysts can be related in a natural way to a $(G,\mathcal{U})$-symmetric catalyst? In general, it is an important problem to constrain potential catalysts at a more basic level than we have done here. In this direction, Theorem 2 gives constraints on potential catalysts that lie beyond the framework of Theorem 1.
@@
Another future direction involving fermions is to make a connection to the notion of fragile topology Po et al. (2018), wherein a topologically nontrivial energy band can be trivialized by the addition of trivial degrees of freedom. This represents an analog of our results for non-interacting fermions. Whether ideas introduced in our paper, possibly in combination with ideas from fragile topology, can be employed to develop an understanding of many body catalysts for interacting fermionic systems would be an interesting topic for future work.
We have also focused on catalyzing short-range entangled phases of matter, and we suspect that catalysts will play a much different role when considering long-range entangled phases. As a preliminary comment, we remark that some of the aspects of our story are reminiscent of the idea of foliated fracton phases of matter Shirley et al. (2018). Therein, certain 3D phases of matter called fracton phases have the property that an FDQC acting on the fracton state and a product state can turn that product state into a topologically ordered state. This process changes the 3D state by reducing its size in one direction, but preserves its phase of matter. So one can view a fracton phase as a catalyst for topologically ordered states, with the caveat that it can only be used finitely many times (assuming a finite system size).
In a similar vein, catalysts can also be used to transform between gapless phases of matter Scaffidi et al. (2017); Verresen et al. (2021). For example, it is shown in Ref. Verresen et al. (2021) that two gapless topological phases with distinct patterns of symmetry enrichment can be related by applying symmetric QCA corresponding to SPT entanglers. Since the catalysts we construct allow such QCA to be implemented on arbitrary states, they can also catalyze the transformation between certain symmetry-enriched gapless phases.